nolegirl24:

Baxter is thinking about buying a car for $18,500. The table below shows the projected value of two different cars for three years. Number of years 1 2 3 Car 1 (value in dollars) 17,390 16,346.60 15,365.80 Car 2 (value in dollars) 17,500 16,500 15,500 Part A: What type of function, linear or exponential, can be used to describe the value of each of the cars after a fixed number of years? Explain your answer. (2 points) Part B: Write one function for each car to describe the value of the car f(x), in dollars, after x years. (4 points) Part C: Baxter wants to purchase a car that would have the greatest value in 9 years. Will there be any significant difference in the value of either car after 9 years? Explain your answer, and show the value of each car after 9 years. (4 points)

nolegirl24:

@Vocaloid PLease help!!!!!!!!!!!!!!!!!!!!!!!!!!

nolegirl24:

So then A is linear right

Vocaloid:

hm. so B is clearly linear but I just double checked A and A isn't linear

Vocaloid:

since the differences in A aren't fixed

nolegirl24:

So the answer to part a is not linear

Vocaloid:

Car 1 is nonlinear (exponential) and Car 2 is linear

nolegirl24:

So then it is exponential

Vocaloid:

make sure to specify which car is which.

nolegirl24:

Okay whats next

Vocaloid:

car 2 is easier to write an equation for it's: value = (rate of change)x + initial value

Vocaloid:

replace "rate of change" with the amount it decreases each year, and "intial value" with 18500

nolegirl24:

We on part b right?

Vocaloid:

yes

nolegirl24:

okay

nolegirl24:

I read what you typed but im confused again.

Vocaloid:

how much does car 2 decrease per year?

nolegirl24:

1,000

Vocaloid:

good take f(x) = (rate of change)x + initial value replace "rate of change" with -1000 and "initial value" with 18500 don't do anything else, let me know what you get

nolegirl24:

Done!

Vocaloid:

what did you get?

nolegirl24:

f(x) = (-1000)x + 18500

Vocaloid:

good, that's car 2 car 1 is a little harder so bear with me

nolegirl24:

Okay awesome! I will

Vocaloid:

if we take the year 1/year 0 value, year 2/year 1 value, year 3/year 2 value, we can see that the car value decreases by 6% each year

Vocaloid:

in other words: value = (value of the previous year)*(0.94)

Vocaloid:

so we can express this as: f(x) = initial value(0.94)^x

nolegirl24:

Wait what is the initial value

Vocaloid:

same as before

Vocaloid:

18500

nolegirl24:

okay

Vocaloid:

f(x) = 18500(0.94)^x is your eq. for car 2 and that's it for b

nolegirl24:

OKyy

Vocaloid:

for part c, take both functions and calculate f(x) when x = 9

Vocaloid:

car 1 f(x) = (-1000)x + 18500 car 2 f(x) = 18500(0.94)^x let x = 9 and find car 1 and car 2

nolegirl24:

So replace x with 9?

Vocaloid:

yes

nolegirl24:

Okay hold on

nolegirl24:

What is the site you used for math???

Vocaloid:

I mean, I used wolframalpha but any online calculator could do fine

nolegirl24:

I got what i needed lol. Hold on im almost done

nolegirl24:

For car 1 is it f(9) = 10600.40384122

Vocaloid:

good, round that to the nearest hundreth since we're talking about money

Vocaloid:

so $10600.40 for car 1 what about car 2?

nolegirl24:

Okay hold on

nolegirl24:

f(9) = 9500

Vocaloid:

good so to answer the question, car 1 is worth more than car 2 by how much?

nolegirl24:

1100.4

Vocaloid:

good, so there ~is~ a significant difference and Baxter should buy car 1 that's it for c

nolegirl24:

Okay

nolegirl24:

Thank you!!